ISS Daily Summary Report – 01/19/16

Electrostatic Levitation Furnace (ELF) Setup:  Kopra and Peake completed the setup of the Japan Aerospace Exploration Agency (JAXA) ELF equipment and installation into the Multi-purpose Small Payload Rack 2 (MSPR2) work volume in the Japanese Experiment Module (JEM).  Later this week crewmembers will continue the setup and next week will install a sample cartridge and checkout the equipment.  The ELF is an experimental facility designed to levitate, melt and solidify materials employing containerless processing techniques that use the electrostatic levitation method with charged samples and electrodes. With this facility, thermophysical properties of high temperature melts can be measured and solidification from deeply undercooled melts can be achieved.

SOLAR: Measurements continued to be taken for European Space Agency’s (ESA’s) SOLAR investigation during the current sun visibility window, which is open from January 11th to January 22nd.  The goal of the SOLAR instruments is to measure solar spectral irradiance and variability.

Nitrogen/Oxygen Recharge System (NORS) Setup and Oxygen Transfer: Kopra set up and configured a NORS Oxygen Recharge Tank and initiated the first transfer of oxygen to the US Airlock High Pressure Gas Tanks (HPGTs) using this new capability.  NORS provides the capability to refill the HPGTs with nitrogen and oxygen following Shuttle retirement. 

Post-Extravehicular Activity (EVA) Activities:  Today, Kopra configured the US Airlock following EVA operations and prepared Extravehicular Mobility Units (EMUs) and equipment for stowage. 

Crew Quarters Baseplate Ballast Assembly (BBA) Replacement:  Kopra replaced the BBA in the starboard Crew Quarters that failed on Sunday night.  The BBA is a component of the General Luminaire Assembly (GLA), which provides illumination inside the Crew Quarters.  The replacement restored lighting to the Crew Quarters.

Node 3 Carbon Dioxide Removal Assembly (CDRA) Fault:  Overnight, the Node 3 (N3) CDRA indicated a fault.  Preliminary data review suggested a Fan Motor Controller (FMC) to Hub Control Zone (HCZ) Multiplexer/Demultiplexer (MDM) communication issue.  All other CDRA hardware data was nominal up to that point.  During the day, the N3 CDRA FMC did not respond to multiple attempts to restart.  The Lab CDRA and the Russian carbon dioxide removal equipment (Vozdukh) are currently up and running.  ISS is in a good configuration for removing carbon dioxide.  Ground teams continue to assess the fault signatures.

Today’s Planned Activities
All activities were completed unless otherwise noted.

  • Laptop RS1(2) Reboot
  • SLEEP – Questionnaire
  • SM Caution & Warning Panel (ПСС) Test
  • [ВКС] Laptops Antivirus software checkout and report
  • PILOT-T. Experiment Prep
  • Video Footage for Nauka 2.0 TV Channel
  • Fine Motor Skills. Examination Ops
  • ELF Components Gathering
  • XF305 Camcorder Setup
  • MSPR2 Work Bench Setup
  • PILOT-T. Experiment Ops.
  • ELF Camcorder Setup
  • PILOT-T. Closeout Ops
  • ИП-1 Flow Sensor Position Check
  • СОЖ Maintenance
  • РСПИ (Data Transmission Radio Link)
  • DAN. Experiment Ops
  • MSPR Ops
  • HD Camcorder 2 Activation in HD Mode
  • Ice Bricks Restow to Original Location in Double Coldbag
  • Hardware Setup for a PAO Event
  • Atmosphere Control and Supply (ACS). Nitrogen and Oxygen Resupply System (NORS). Oxygen Transfer
  • IPAD Reconfig for Dose Tracker
  • Water Recovery Management (WRM). Condensate Pumping (start)
  • HABIT. Documentation of the Examination Results.
  • EVA Hardware and Tools Search
  • ARED. Cylinder Flywheel Evacuation
  • Preparation of materials for Roscosmos press center re EVA-42 [Aborted – to be rescheduled to allow RS additional crewmember participation]
  • Switching the HD Camcorder 2 to SD mode
  • HABIT. Questionnaire
  • Water Recovery Management (WRM). Condensate Pumping
  • Hardware Transfer to Airlock after US EVA
  • Delta File Prep
  • CONTENT. Experiment Ops
  • Checkout of P/TV camcorder setup
  • HMS. Nutritional Assessment (ESA)
  • Still Cameras Reconfig for EVA
  • TWIN. Urine Sample Collection
  • Regeneration of EMU purification cartridge
  • BSA Battery Maintenance
  • ARED Exercise Video Hardware Restow
  • VZAIMODEISTVIE-2. Experiment Ops

Completed Task List Items

  • None

 Ground Activities

All activities were completed unless otherwise noted.

  • Configuring ISS for NORS operations

Three-Day Look Ahead:

Wednesday, 01/20: EMU Loop Scrubs, Airway Monitoring, Cardox, Twin Study
Thursday, 01/21: ISS Emergency Training, Fundoscope, Airway Monitoring
Friday, 01/22: RRM Taskboard 4 Removal from JEMAL, ELF, OcularHealth, HMS Ultrasounds, Airway Monitoring Stow







[СКВ] 1 – SM Air Conditioner System (“SKV1”)


[СКВ] 2 – SM Air Conditioner System (“SKV2”)


Carbon Dioxide Removal Assembly (CDRA) Lab


Carbon Dioxide Removal Assembly (CDRA) Node 3


Major Constituent Analyzer (MCA) Lab


Major Constituent Analyzer (MCA) Node 3


Oxygen Generation Assembly (OGA)


Urine Processing Assembly (UPA)


Trace Contaminant Control System (TCCS) Lab

Full Up

Trace Contaminant Control System (TCCS) Node 3